OpenStudy (anonymous):

solve by the substitution method 5x + 5y = 20, -3x + y = 12

OpenStudy (anonymous):

-3x + y = 12 y = 12 + 3x substitue y = 12 + 3x into 5x + 5y = 20 and solve

OpenStudy (anonymous):

? not sure

OpenStudy (anonymous):

not sure about what

OpenStudy (anonymous):

you typed sub y = 12 + 3x into 5x + 5y =20 what am i substituting?

OpenStudy (anonymous):

you know that y equals 12 + 3x In this equation there's a y 5x + 5y =20

OpenStudy (anonymous):

got it. thanks